# Revisiting Intermediate-Layer Matching in Knowledge Distillation: Layer-Selection Strategy Doesn’t Matter (Much)

### v1

@@ 3.2 Datasets and Models
Figure 1: (a) Illustration of the angle calculation. Cosine similarities are shown for (b) MNLI classification, (c) Encoder in the WMT task, and (d) Decoder in WMT. Orange refers to the setup of random parameter initialization and blue refers to student weights initialized by the teacher.
@@
For the student, we adopted the teacher’s architecture but reduced the number of layers to three. Note that, for DART and WMT16, we had three layers for the encoder and another three layers for the decoder. Moreover, we employed two parameter initialization strategies for the student: randomly initializing the weights and copying the weights from the corresponding teacher layer. The former isolates the effects of intermediate-layer matching from weight copying, whereas the latter is a more practical method that yields higher performance Sanh, (2019); Shleifer and Rush, (2020).
4 Results and Analysis
Main Results. In Table 1, we present the main results of our layer-selection experiments. As seen in Lines 1–2 , our finetuned teacher models perform similarly to previous work Devlin et al., (2019); Nan et al., (2021); Wen et al., (2023), showing that we have successfully set up the environment for KD experiments.
We examine the performance of different layer-selection strategies. As shown in Lines 4–7 and 9–12, the student model achieves similar results across different strategies, with only 2–3 points difference in accuracy for classification tasks and 1–2 points difference in BLEU for generation tasks. Notice that Reverse Matching and Random Matching appear nonsensical, when in fact they still achieve close performance to Forward Matching, often outperforming No Matching. The results show that layer-selection strategy has an unexpectedly small effect on student performance; this highlights the limitations of previous research on layer-selection strategies.
It should be emphasized that intermediate-layer matching indeed helps KD compared with No Matching [^3] One exception is the DART experiment with randomly initialized weights, for which we suspect intermediate-layer matching causes the student to overfit. That said, different strategies still perform similarly to conventional Forward Matching, and thus, it does not contradict our general finding., even though the matching strategy does not play a significant role. On MNLI, for example, all strategies improve upon No Matching by ten points in the setting of random initialization and two points when the student weights are initialized from the teacher.
Next, we take a closer look at how different layer-selection strategies behave under the two parameter initialization settings. To reiterate, copying the teacher’s parameters for initialization is a simple and practical method to quickly transfer the teacher’s knowledge to the student Sanh, (2019); Shleifer and Rush, (2020). In our experiments, it is evident that parameter copying indeed leads to significant improvements compared to random initialization. Nonetheless, the general trend is consistent: intermediate-layer matching is important, while layer-selection methods do not matter (much).
The Angles of Matching Different Layers. A curious question arises from these observations: why does intermediate-layer matching help KD but different layer-selection strategies perform similarly? To answer this, we measure the angles between the teacher’s layers, viewed from the student. Specifically, we measure the angles formed by two teacher layers’ and one student layer’s vector representations, depicted in Figure 1a. We show the phenomenon in the MNLI and WMT16 En–Ro datasets in Figure 1b, 1c and 1d. We see that in both randomly initialized and weight-copied settings, the cosine similarity is positive, suggesting that the angles are mostly acute. In other words, the student layer is pulled to the same general direction regardless of which teacher layer it is matched to. This explains the findings in our paper.
Appendix. We provide additional analysis on the student depth in Appendix A.
5 Conclusion
In this paper, we observe an intriguing phenomenon that although intermediate-layer matching helps knowledge distillation, the layer-selection strategy does not matter (much); we also provide an interpretation based on the angles of teacher and student layers. Our work suggests potential limitations and oversights in previous work, where researchers present various heuristic layer matching methods when training their distilled systems, but their effect is not comprehensively studied.
6 Limitations
In our work, we have experimented with various setups, including six tasks (four classification and two generation), three model architectures, and two parameter initialization methods. Although the results are generally consistent, there is one exception that intermediate-layer matching does not help in the DART setup. Nevertheless, this is understandable as empirical findings are usually not theoretically guaranteed.
It is also worth mentioning that our work does not suggest intermediate-layer matching is unhelpful for KD. Rather, we present an interesting phenomenon that the layer-selection strategy plays an insignificant role in the process. We argue that future studies on layer selection should have closer examination and more rigorous comparison on its effect.
Appendix A Analysis of Student Depths

### v2

@@ Appendix A Datasets and Models
CommonsenseQA. The CommonsenseQA dataset (commonsense) is a question-answering task where each question is followed by five possible answers. We concatenated the correct answer with the question, then finetuned the 40-layer Qwen3-8B for next-token prediction to use as the teacher. Following lm_eval, we use zero-shot prompting to obtain responses from the language model and report exact match with the ground truth.
HellaSwag. The HellaSwag dataset (hellaswag) is a difficult sentence completion task that presents prefix sequence along with four possible endings. Like the previous task, we finetuned a 40-layer Qwen3-8B model for next-token prediction. We follow lm_eval and use the language model to rank each ending according to perplexity, conditioned on the prefix.
@@
For the student, we adopted the teacher’s architecture but reduced the number of layers to three in the main experiments. Specifically, we use teacher Layers 4, 8, and 12 for matching. Note that, for BART and T5 models, this means three layers for the encoder and another three layers for the decoder. Moreover, we employed two parameter initialization strategies for the student: randomly initializing the weights and copying the weights from the corresponding teacher layer. The former isolates the effects of intermediate-layer matching from weight copying, whereas the latter is a more practical method that yields higher performance distilBERT; predistill.
Regarding the more challenging experiments, namely HellaSwag and CommonsenseQA, we shrunk the student by the same proportions as in the main setup. Specifically, we distilled the 40-layer teacher to a 10-layer student, using Layers 4, 8, $\cdots$, and 40 in the teacher model for matching.
Appendix B Analysis of Student Depths
@@
Appendix C Experimental Stability
In the main results (Table 1), excluding the Out-of-Order Random Matching setup, we ran every experiment only once due to the large number of models, tasks, and setups.
We show the stability of our results in Table 4 by computing the standard deviation of five runs. Here, we chose Out-of-Order Random Matching because this is in theory the most noisy setup due to the stochasticity of layer matching.
In Table 4, we see that random initialization yields higher standard deviation than the weight-copied setting. This is understandable, as the former setup involves more randomness. Nevertheless, the model performs stably in both settings, showing that our results and findings are reliable.